OpenStudy (anonymous):

PLEASE HELP ME! The point (-2,k) lies on the circle x2+y2=20. Determine the values of k. Show all steps in the solution.

hero (hero):

Plug (x,y) = (-2,k) in to the equation x^2 + y^2 = 20 \((-2)^2 + (k)^2 = 20\) \(4 + k^2 = 20\) \(k^2 = 16\) \(k = ±4\)

OpenStudy (anonymous):

how did u get 16?

hero (hero):

by subtracting 4 from both sides

OpenStudy (anonymous):

and that gives u 4 in the end then.

OpenStudy (anonymous):

how did u get 4 in the end?

hero (hero):

The steps are posted above